OpenStudy (anonymous):

determine u and v are orthogonal, parallel, or neither. u=(15, 45) v=(-5, 12)

OpenStudy (anonymous):

take the dot product, if the dot product is zero, they are orthogonal the dot product here is 15(-5) + 45(12) so the vectors are not orthogonal to see if they are parallel, take the dot product again, using the definition u dot v = mag u x mag v cos (theta) where mag is the magnitude of the vector and theta is the angle between them we know from above that u dot v = 465 mag u = sqrt[15^2+45^2]=47.43 mag v = sqrt[5^2+12^2]=13 so we know: cos(theta) = 465/(47.43x13) = 0.75 so theta = 41 deg

OpenStudy (anonymous):

thanks very much, just wanted to make sure it was neither

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

To see that they are not parallel, notice that u is not equal a scalar v if u=(15, 45) v=(-5, 12) \[ u= a ( v)\\ 15 = - 5 a \\ 45 = 12 a \\ a= -3\\ a =\frac { 45}{12}\ne -3 \]